5.00 mL of 0.250 M ammonia (NH3) is titrated with 0.100 M hydrochloric acid (HCl). The...

  1. 5.00 mL of 0.250 M ammonia (NH3) is titrated with 0.100 M hydrochloric acid (HCl). The Kb for ammonia is 1.75 x 10-5
    What is the pH of the ammonia solution prior to the addition of any HCl?
    What is the pH of the solution after 5.00 mL of HCl has been added?
Answer #1

initially only NH3 present.

for weak acid

pOH = 1/2 [pKb - log C]

pKb = - log Kb = - log [1.75 x 10-5] = 4.76

pOH = 1/2 [4.76 - log 0.250]

pOH = 2.68

pH = 14 - 2.68

pH = 11.32

2) initially

millimoles of NH3 = 5.00 x 0.250 = 1.250

millimoles of HCl added = 5.00 x 0.100 = 0.50

after HCl added

millimoles of NH3 left = 1.250 - 0.50 =0.750

millimoles of NH4+ = 0.50

total volume = 5.00 + 5.00 = 10.00 mL

[NH3] = 0.750 / 10.00 = 0.075 M

[NH4+] = 0.50 / 10.00 = 0.05 M

pOH = pKb + log [NH4+] / [NH3]

pOH = 4.76 + log [0.05] / [0.075]

pOH = 4.58

pH = 14 - 4.58

pH = 9.42